Solve the triangle ABC given that A=66 degrees,C=44 degrees and a = 7cm.


(Hint: find the other sides and the area)


1
Expert's answer
2021-09-17T03:54:27-0400

Given that "\angle A=66\degree,\angle C=44\degree,a=7 cm"

"\therefore \angle B=70\degree"

Now using sine law, we get:

"\frac{sinA}{a}=\frac{sinB}{b}=\frac{sinC}{c}\\
\Rightarrow \frac{sin66\degree}{7}=\frac{sin70\degree}{b}=\frac{sin44\degree}{c}\\
\Rightarrow \frac{0.91}{7}=\frac{0.93}{b}=\frac{0.69}{c}\\
\therefore b=7.15cm;\ c=5.3 cm"

Now, area of "\triangle ABC=\frac{1}{2}bcsinA\\"

"=\frac{1}{2}\times (7.15)\times (5.3)\times (sin66\degree)\\
=\frac{1}{2}\times (7.15)\times (5.3)\times (0.91)\\
=17.24 \ cm^2"